Campbell Supper Co. paid a $0.662 dividend per share in 2013, which grew to $0.80 in 2016. This growth is expected to continue. What is the value of this stock at the beginning of 2017 when the required return is 8.3 percent?
Answer #1

FV = PV(1+g)^n

0.80 = 0.662 ( 1+ g)^3

(1+g)^3 = 0.8 / 0.662

= 1.2085

1+g = 1.2085 ^ ( 1/3)

= 1.0651

g = 1.0651 - 1

= 0.0651 i.e 6.51%

P0 = D1 / [ Ke - g ]

D1 = D0 (1+g)

= 0.80 ( 1 + 0.0651 )

= 0.80 (1.0651)

= 0.8521

P0 = D1 / [ Ke - g ]

= 0.8521 / [ 8.3% - 6.51% ]

= 0.8521 / 1.79%

= $ 47.60